OpenStudy (anguyennn):

Hello could I get some help with evaluating 2√5/√7

OpenStudy (kittiwitti1):

\[2\sqrt{5}\sqrt{7}\text{ or }2\sqrt{5\sqrt{7}}?\]

OpenStudy (anguyennn):

none of those

OpenStudy (anguyennn):

here ill draw it out for you

OpenStudy (anguyennn):



OpenStudy (anguyennn):

I want to know how to do it

OpenStudy (anguyennn):

algebraically

OpenStudy (triciaal):